Columbia Art Studio sells framed art prints for $100. The unit variable cost per print is $50 plus a selling commission of 10%. Fixed manufacturing costs total $1,250 per month, while fixed selling and administrative costs total $2,500.

How many prints must be sold to earn pretax income of $7,500?
Answer #1

Selling price per unit = $100

Variable cost per unit = $50+ 10% sales commission

= 50 + 100 x 10%

= $60

Fixed manufacturing costs = $1,250

Fixed selling and administrative costs = $2,500

Total fixed costs = Fixed manufacturing costs + Fixed selling and administrative costs

= 1,250+2,500

= $3,750

Contribution margin per unit= Selling price per unit - Variable cost per unit

= 100-60

= $40

Number of units to be sold for target profit = ( Fixed cost + Target profit ) / Contribution margin per unit

= (3,750+7,500)/40

= 11,250/40

= 282 units (281 (if rounded to nearest whole number)

Kindly comment if you need further assistance. Thanks